# Key Words.: cosmology: distance scale – gravitational lensing: strong – Galaxy: kinematics and dynamics – Galaxies: elliptical and lenticular, cD1 Introduction

### v1

@@
Stellar velocity dispersion () of massive elliptical galaxies is a key ingredient to breaking the mass-sheet degeneracy and obtaining precise and accurate cosmography from gravitational time delays. The relative uncertainty on the Hubble constant H0 is double the relative error on . Therefore, time-delay cosmography imposes much more demanding requirements on the precision and accuracy of than galaxy studies. While precision can be achieved with an adequate signal-to-noise ratio (SNR), accuracy depends critically on factors such as elemental abundances and temperature of stellar templates, flux calibration, and wavelength ranges. We carry out a detailed study of the problem using multiple sets of galaxy spectra of massive elliptical galaxies with SNR30–160 Å-1, and state-of-the-art empirical and semi-empirical stellar libraries and stellar population synthesis templates. We show that the choice of stellar library is generally the dominant source of residual systematic error. We propose a general recipe to mitigate and account for residual uncertainties. We show that sub-percent accuracy can be achieved on individual spectra with our data quality. Covariance between velocity dispersions measured for a sample of spectra can also be reduced to sub-percent levels. We recommend this recipe for all applications that require high precision and accurate stellar kinematics, and make all software publicly available to facilitate its implementation. This recipe will be used in future TDCOSMO collaboration papers.
@@
Stellar velocity dispersion () has been a key observable for our understanding of massive elliptical galaxies for decades (Faber & Jackson 1976). Advances in spectroscopy, stellar populations and dynamical models have enabled the understanding of their internal structure, orbital anisotropy, mass-to-light ratio, and many other fundamental properties. The tight correlations between stellar velocity dispersion and other parameters are a testbed for any galaxy formation model (see review by Cappellari 2016).
@@
In the past two decades, it has been shown that stellar velocity dispersion is particularly powerful in combination with strong gravitational lensing (e.g., Treu 2010, and references therein). As a purely dynamical probe, stellar velocity dispersion probes the full three-dimensional structure of galaxies and is insensitive to the mass-sheet degeneracy (Koopmans et al. 2003). Conversely, lensing probes the total projected mass (Dutton et al. 2011) and is insensitive to the mass anisotropy degeneracy (Mamon & Łokas 2005; Courteau et al. 2014).
@@
For all these reasons, stellar velocity dispersion measurements have assumed a central role in the effort to infer cosmological parameters using gravitational time delays in galaxy-scale lenses, hereafter time delay cosmography (Suyu et al. 2010; Treu & Marshall 2016; Treu et al. 2022; Birrer et al. 2024). By constraining the mass distribution of the deflector (typically a massive elliptical galaxy) (Treu & Koopmans 2002) and by breaking the mass-sheet degeneracy, stellar velocity dispersions can, in principle, help achieve percent-level precision on H0, particularly when spatially resolved kinematics is available (Shajib et al. 2018; Yıldırım et al. 2020; Birrer & Treu 2021).
@@
However, in order for velocity dispersions to be a useful ingredient for time delay cosmography, precision and accuracy of better than 1% are needed, since the relative error on H0 scales as double the error on stellar velocity dispersion: (e.g. Chen et al. 2021). Precision is relatively easy to achieve with modern instruments: spectra with current SNR and resolution are more than sufficient to obtain formal errors of order 1%. Furthermore, if errors are random and uncorrelated, it is sufficient to average multiple systems and/or measurements to obtain the desired precision. In contrast, potential systematic errors may bias individual measurements or introduce covariance between different galaxies – or within parts of the same galaxy – thus introducing an effective systematic noise floor, independent of the SNR or number of systems in the sample. Sources of systematics in stellar velocity dispersion measurements have been discussed for decades, and include instrumental resolution, bias due to insufficient signal-to-noise ratio, wavelength range, mismatch between templates and galaxy spectra in terms of elemental abundances and temperature, mismatch between template and galaxy continua, imperfect flux calibration, and interstellar absorption. Many excellent studies on this topic have been published in the past and inform the work presented here (e.g., Dressler 1984; Kelson et al. 2000; Treu et al. 2001; Barth et al. 2002; Spiniello et al. 2015, 2021a, 2021b; Mehrgan et al. 2023; D’Ago et al. 2023).
@@
In this paper, we revisit the issue of systematic errors in measured stellar velocity dispersion using multiple sets of galaxy spectra with SNR 30-160 Å-1 and state-of-the-art templates and fitting code. Our goal is twofold: minimize systematic errors and quantify residual systematics in a principled, general, and reproducible way. In order to achieve our goal, we proceed as follows. In Section 2, we assemble a high-quality set of stellar and stellar population synthesis templates. The high-quality set of templates is made public for use by third parties. In Section 3, we present sets of spectra of elliptical galaxies at observed with four different instruments (MUSE on VLT, KCWI on Keck, JWST-NIRSpec, SDSS; Bacon et al. 2010; Morrissey et al. 2018; Jakobsen et al. 2022). The SDSS spectra are included to compare with previous studies. In contrast, the three other sets of spectra are meant to represent the current state-of-the-art instruments used for stellar kinematics of lens galaxies.
@@
In Section 4, we examine, quantify, and mitigate all known sources of systematics by running comprehensive sets of measurements. For all our fits, we use the well-tested and established code pPXF111https://pypi.org/project/ppxf/ (Cappellari & Emsellem 2004; Cappellari 2017, 2023). To guard against bugs in the utilization of pPXF and the construction of stellar libraries, SK and PM developed two independent pipelines. Several tests show that the two pipelines yield identical results when applied to the same data.
@@
In Section 5, we propose a procedure to account for residual systematics by treating unconstrained choices (chiefly the choice of the stellar library) as nuisance parameters and marginalizing over them. The procedure can be used to estimate random and systematic errors as well as covariance between sets of measurements, e.g., spatial bins within a galaxy.
@@
In Section 6, we show that by applying our procedure to high-quality spectra of elliptical galaxies, both systematic errors and covariance between independent galaxies or spatial bins can be contained to a sub-percent level. Our specific results only apply to the conditions described in this paper, in terms of instrumental setup, template libraries, and type of spectra under consideration. However, our procedure is entirely general and can be applied to any set of measurements to estimate the corresponding systematic noise floor. We recommend our recipe for all applications that require high precision and accurate stellar kinematics, and we make all software publicly available to facilitate its implementation. This recipe will be used in future TDCOSMO collaboration papers.
In Section 7, we summarize and discuss our results.
2 Stellar libraries and Single Stellar Population (SSP) models
@@
In this section, we briefly describe the stellar libraries used in this work, referring to the original papers for more details (§ 2.1). We then describe our rigorous process to “clean” the libraries in Section 2.2. In Section 2.3, we compare the parameter space covered by the libraries relative to each other and before and after cleaning, as well as the spectra of the stars that are in common between the libraries.
@@ 2.1 Description of libraries
We consider two types of templates for the kinematic extraction: (i) individual stellar spectra and (ii) stellar-population synthesis models based on those stellar libraries. As empirical stellar libraries, we use three that are commonly employed in stellar kinematics determination: Indo-US222https://noirlab.edu/science/observing-noirlab/observing-kitt-peak/telescope-and-instrument-documentation/cflib (Valdes et al. 2004), MILES333https://miles.iac.es/ (Sánchez-Blázquez et al. 2006; Falcón-Barroso et al. 2011), and X-shooter Spectral Library (XSL)444http://xsl.u-strasbg.fr/ (Verro et al. 2022). Additionally, we include the sMILES stellar library, a semi-empirical library derived by modifying the MILES stellar spectra to represent different element abundances (Knowles et al. 2021). This modification is meant to address template mismatch issues often encountered in the study of massive elliptical galaxies (Barth et al. 2002; Mehrgan et al. 2023). However, it is important to note that these models have known limitations, particularly in the blue part of the optical spectrum (Knowles et al. 2021, 2023). Therefore, our expectation is that they would not be sufficiently accurate for 1% kinematics. We include them for completeness.
As stellar population synthesis models, also known as single stellar populations (SSP), we use the E-MILES and sSMILES, (hereafter E-MILES SSP and sMILES SSP, Vazdekis et al. 2010, 2016; Knowles et al. 2023) and the SSP based on the XSL library (hereafter XSL SSP, Verro et al. 2022). While these models offer a physically motivated mix of stars, they lack the flexibility provided by individual stellar spectra. Therefore, they cannot reproduce the complexity of real star formation histories and chemical enrichment histories and are unlikely to be sufficient for sub-percent accuracy. However, they have been popular choices in the past. Therefore, we investigate them for completeness in order to determine how well they agree with the stellar libraries.
The Indo-US and MILES libraries have an approximately constant full width at half maximum (FWHM) kinematic resolution of 1.36 Å and 2.5 Å, respectively (Falcón-Barroso et al. 2011), which correspond to 43 km s-1 and 80 km s-1, respectively, at 4000 Å. For all MILES-based libraries (E-MILES, sMILES, etc.), we adopted 2.5 Å as the FWHM resolution of the templates. The kinematic resolution of the XSL templates is km s-1 at the UVB region (3000–5560 Å), 11 km s-1 at the VIS region (5330–10200 Å), and 16 km s-1 at the NIR region (9940–24800 Å). We refer the reader to the original papers for a full description of the libraries.
@@ 2.2 Cleaning the stellar libraries
The full stellar libraries contain spectra of variable quality, including some with defects, incomplete data, incomplete calibration, or low SNR. Indeed, when using stellar spectra to construct stellar population synthesis models, the original papers apply quality cuts (Vazdekis et al. 2016; Verro et al. 2022).
In order to obtain “clean” spectral libraries, we took a series of steps. First, we eliminated all the spectra that were flagged by the original teams as faulty. For example, for XSL, the template FITS file headers contain keywords that indicate the applied corrections and quality flags, and some of the spectra were flagged as “peculiar”, “abnormal”, “zero” (see, e.g., Tables B.1 and C.1 of Verro et al. 2022). In addition, some spectra have not been corrected for slit losses. We eliminated from XSL all spectra listed in the aforementioned Table C.1 and all that were corrected for slit loss. Whenever possible we contacted the authors of the library to reproduce the selection applied to their construction of the stellar population synthesis models.
Second, we eliminated all the spectra with incomplete stellar parameters, especially temperature Teff and metallicity Fe/H. For MILES and XSL, they represent a small fraction (1-2%) of the total number of stars, while for Indo-US, they are a larger fraction (20%). In almost all cases, they are faulty spectra (e.g., missing part of the spectral range, very noisy, or with obvious defects). We note that many of the spectra flagged by the authors of the MILES and XSL libraries are also missing and Fe/H. These two initial steps remove between 20-30% of the initial spectra for all three libraries. For reference, the Indo-US library has been used in the most recent versions of the SDSS kinematic pipeline (Bolton et al. 2008, and Bolton, A.S. private communication). However, to our knowledge, no cleaning was applied.
Third, we visually inspected all the individual spectra to identify defects or anomalies. Although, in general, the quality of the spectra is excellent, in very few cases, there were substantial anomalies (see Table 1 and Figure 1). The spectra were flagged for the following anomalies:
@@
Improper flux calibration at the edges and near the dichroic. These spectra were kept in the clean libraries. However, we recommend avoiding/masking the regions near the edge and near the dichroic when performing the fits.
@@
Some spectra, more commonly from the Indo-US library, have localized gaps that can extend for 100s or 1000s of Å. We do not exclude them from the clean library, but we recommend excluding them if the gaps overlap with the fitted range.
@@
After the cleaning procedure, the stellar libraries consist of 789, 496, and 992 stars, as summarized in Table 1. The list of stars in the “clean” stellar libraries will be made available in electronic format through GitHub https://github.com/TDCOSMO/KINEMATICS_METHODS after the acceptance of this manuscript.
@@
After library authors’ cuts | 801 | 530 | 1273 |
With and Fe/H | 790 | 522 | 1006 |
After visual Inspection | 789 | 496 | 992 |
2.3 Comparing parameter space and common stellar spectra
In order to get an unbiased estimate of the stellar velocity dispersion, it is important to use stellar libraries that span the range of temperature and metallicity expected to dominate the spectra of massive elliptical galaxies. The distribution of these two parameters for the stellar libraries before and after cleaning is unaffected, demonstrating that the quality cuts do not exclude important parts of the parameter space and remain representative of the original libraries.
We also tested whether the distribution of the stellar libraries within the parameter space affects the outcome. Figure 2 shows that with spectra of sufficient quality and wavelength coverage, pPXF is able to recover the correct stellar velocity dispersion and stellar mix, irrespective of the distribution in temperature and metallicity of the stellar spectra within the library.
For the clean stars in common between the libraries, we compared the spectra visually and in terms of their ratio. Some examples are shown in Figure 3. Overall, the comparison is excellent, even though there are some instances in which the ratio differs significantly from one, indicating systematic residuals in the flux calibration and reddening corrections. These differences underscore the importance of having additive and multiplicative polynomial continuum components in the kinematic fits, as discussed below.

### v2

# with sub-percent accuracy for cosmography

@@
with sub-percent accuracy for cosmography
The stellar velocity dispersion () of massive elliptical galaxies is a key ingredient in breaking the mass-sheet degeneracy and obtaining precise and accurate cosmography from gravitational time delays. The relative uncertainty on the Hubble constant H0 is double the relative error on . Therefore, time-delay cosmography imposes much more demanding requirements on the precision and accuracy of than galaxy studies. While precision can be achieved with an adequate signal-to-noise ratio (S/N), the accuracy critically depends on key factors such as the elemental abundance and temperature of stellar templates, flux calibration, and wavelength ranges. We carried out a detailed study of the problem using multiple sets of galaxy spectra of massive elliptical galaxies with S/N30–160 Å-1, along with state-of-the-art empirical and semi-empirical stellar libraries and stellar population synthesis templates. We show that the choice of stellar library is generally the dominant source of residual systematic errors. We propose a general recipe for mitigating and accounting for residual uncertainties. We show that a sub-percent level of accuracy can be achieved on individual spectra with our data quality, which we subsequently validated with simulated mock datasets. The covariance between velocity dispersions measured for a sample of spectra can also be reduced to sub-percent levels. We recommend this recipe for all applications that require high precision and accurate stellar kinematics. Thus, we have made all the software publicly available to facilitate its implementation. This recipe will also be used in future TDCOSMO collaboration papers.
@@
The stellar velocity dispersion () has been a key observable for our understanding of massive elliptical galaxies for decades (Faber76). Advances in spectroscopy, stellar populations, and dynamical models have heightened our understanding of their internal structure, orbital anisotropy, mass-to-light ratio, and many other fundamental properties. The tight correlations between stellar velocity dispersion and other parameters serve as a testbed for any galaxy formation model (see a review by Cappellari16).
@@
Over the past two decades, the stellar velocity dispersion has been shown to be particularly powerful in combination with strong gravitational lensing (e.g., Treu10b, and references therein). As a purely dynamical probe, the stellar velocity dispersion allows us to investigate the full three-dimensional structure of galaxies, while remaining insensitive to the mass-sheet degeneracy (Koopmans03). Conversely, lensing probes the total projected mass (Dutton11) and is insensitive to the mass anisotropy degeneracy (Mamon05; Courteau14).
@@
For all these reasons, stellar velocity dispersion measurements have assumed a central role in the effort to infer cosmological parameters using gravitational time delays in galaxy-scale lenses; hereafter, this is referred to as time-delay cosmography (Suyu10; Treu16b; Treu22; Birrer24). By constraining the mass distribution of the deflector (typically a massive elliptical galaxy) (treu02) and by breaking the mass-sheet degeneracy, stellar velocity dispersions can, in principle, help achieve percent-level precision on H0, particularly when spatially resolved kinematics is available (Shajib18; Yildirim20; B+T21).
@@
However, for velocity dispersions to be a useful ingredient for time-delay cosmography, a level of precision and accuracy better than 1% is needed, since the relative error on H0 scales as double the error on stellar velocity dispersion, expressed as (e.g. Chen21). Precision is relatively easy to achieve with modern instruments, namely: spectra with current S/N and resolution are more than sufficient to obtain formal errors of order 1%. Furthermore, if errors are random and uncorrelated, it is sufficient to average multiple systems and/or measurements to obtain the desired precision. In contrast, potential systematic errors might bias individual measurements or introduce covariance between different galaxies (or within parts of the same galaxy) thereby introducing an effective systematic noise floor, independent of the S/N or number of systems in the sample. Sources of systematics in stellar velocity dispersion measurements have been a topic of discussion for decades, including instrumental resolution, bias due to insufficient S/N, wavelength range, mismatch between templates and galaxy spectra in terms of elemental abundances and temperature, mismatch between template and galaxy continua, imperfect flux calibration, and interstellar absorption. Many excellent studies on this topic have been published in the past, and their findings inform the work presented here (e.g., Dressler84; Kelson00; Treu01; Barth02; Spiniello15; Spiniello21a; Spiniello21b; Mehrgan2023; DAGO23).
@@
In this paper, we revisit the issue of systematic errors in measured stellar velocity dispersion using multiple sets of galaxy spectra with S/N 30-160 Å-1, along with state-of-the-art templates and fitting codes. Our goal is twofold: minimize systematic errors and quantify residual systematics in a principled, generalized, and reproducible way. Our paper is organized as follows. In Section 2, we assemble a high-quality set of stellar and stellar population synthesis templates. The high-quality set of templates has been made public for use by third parties. In Section 3, we present sets of spectra of elliptical galaxies at observed with four different instruments (MUSE on VLT, KCWI on Keck, JWST-NIRSpec, SDSS; MUSE; KCWI; NIRSpec). The SDSS spectra are included for the sake of comparisons with previous studies. In contrast, the three other sets of spectra are meant to represent current state-of-the-art instruments used for stellar kinematics of lens galaxies.
@@
In Section 4, we examine, quantify, and mitigate all known sources of systematics by running comprehensive sets of measurements. For all our fits, we use the well-tested and established code pPXF111https://pypi.org/project/ppxf/ (Cappellari04; Cappellari17; Cappellari23). To guard against bugs in the utilization of pPXF and the construction of stellar libraries, SK and PM developed two independent pipelines. Several tests show that the two pipelines yield identical results when applied to the same data.
@@
In Section 5, we propose a procedure to account for residual systematics by treating unconstrained choices (chiefly the choice of the stellar library) as nuisance parameters and marginalizing over them. The procedure can be used to estimate random and systematic errors as well as covariance between sets of measurements (e.g., spatial bins within a galaxy).
@@
In Section 6, we show that by applying our procedure to high-quality spectra of elliptical galaxies, both systematic errors and covariance between independent galaxies or spatial bins can be contained to a sub-percent level. Our specific results only apply to the conditions described in this paper, in terms of instrumental setup, template libraries, and types of spectra under consideration. However, our procedure is entirely general and can be applied to any set of measurements to estimate the corresponding systematic noise floor. We recommend our recipe for all applications that require high precision and accurate stellar kinematics. We have made all the software publicly available to facilitate its implementation. This recipe will be used in future TDCOSMO collaboration papers. Finally, in Section 7, we summarize and discuss our results.
2 Stellar libraries and single stellar population models
@@
In this section, we briefly describe the stellar libraries used in this work, referring to the original papers for more details (Section 2.1). We then describe our rigorous process to “clean” the libraries in Section 2.2 and compare spectra of the stars in common between the libraries.
@@ 2.1 Description of libraries
We considered two types of templates for the kinematic extraction: (i) individual stellar spectra and (ii) stellar-population synthesis models based on those stellar libraries. As the empirical stellar libraries, we used three that are commonly employed in stellar kinematics determination: Indo-US222https://noirlab.edu/science/observing-noirlab/observing-kitt-peak/telescope-and-instrument-documentation/cflib (Valdes04), MILES333https://miles.iac.es/ (SanchezBlazquez2006; Falcon-Barroso11), and X-shooter Spectral Library (XSL)444http://xsl.u-strasbg.fr/ (Verro22). Additionally, we included the sMILES stellar library, a semi-empirical library that was derived by modifying the MILES stellar spectra to represent different element abundances (Knowles21). This modification is meant to address template mismatch issues often encountered in the study of massive elliptical galaxies (Barth02; Mehrgan2023). However, it is important to note that these models have known limitations, particularly in the blue part of the optical spectrum (Knowles21; Knowles23). Therefore, our expectation is that they would not be sufficiently accurate for 1% kinematics. We included them in our study for completeness.
For our stellar population synthesis models, also known as single stellar population (SSP) models, we used the E-MILES and sSMILES, (hereafter E-MILES SSP and sMILES SSP, Vazdekis10; Vazdekis16; Knowles23) as well as the SSP model based on the XSL library (hereafter XSL SSP, Verro22b). While these models offer a physically motivated mix of stars, they lack the flexibility provided by individual stellar spectra. Therefore, they cannot reproduce the complexity of real star formation histories and chemical enrichment histories and are unlikely to be sufficient for sub-percent accuracy. However, they have been popular choices in the past. Therefore, we investigated them for completeness to determine how well they agree with the stellar libraries.
The Indo-US and MILES libraries have an approximately constant full width at half maximum (FWHM) kinematic resolution of 1.36 Å and 2.5 Å, respectively (Falcon-Barroso11), which correspond to 43 km s-1 and 80 km s-1, respectively, at 4000 Å. For all MILES-based libraries (E-MILES, sMILES, etc.), we adopted 2.5 Å as the FWHM resolution of the templates. The kinematic resolution of the XSL templates is km s-1 at the UVB region (3000–5560 Å), 11 km s-1 at the VIS region (5330–10200 Å), and 16 km s-1 at the NIR region (9940–24800 Å). We refer to the original papers for a full description of the libraries.
@@ 2.2 Cleaning the stellar libraries
The full stellar libraries contain spectra of variable quality, including some with defects, incomplete data, incomplete calibration, or low S/N. Indeed, when using stellar spectra to construct stellar population synthesis models, the original papers applied quality cuts (Vazdekis16; Verro22).
To obtain “clean” spectral libraries, we took a series of steps. First, we eliminated all the spectra that were flagged by the original teams as faulty. For example, for XSL, the template FITS file headers contain keywords that indicate the applied corrections and quality flags and some of the spectra were flagged as “peculiar,” “abnormal,” and “zero” (see, e.g., Tables B.1 and C.1 of Verro22). In addition, some spectra have not been corrected for slit losses. We eliminated from XSL all spectra listed in the aforementioned Table C.1 and all that were corrected for slit loss. Whenever possible, we contacted the authors of the library to reproduce the selection applied to their construction of the stellar population synthesis models.
Second, we eliminated all the spectra with incomplete stellar parameters, especially the temperature, Teff, and metallicity, Fe/H. For MILES and XSL, they represent a small fraction (1-2%) of the total number of stars, while for Indo-US, they are a larger fraction (20%). In almost all cases, they are faulty spectra (e.g., missing part of the spectral range, very noisy, or with obvious defects). We note that many of the spectra flagged by the authors of the MILES and XSL libraries are also missing and Fe/H values. These two initial steps remove between 20-30% of the initial spectra for all three libraries. For reference, the Indo-US library has been used in the most recent versions of the SDSS kinematic pipeline (Bolton08, and Bolton, A.S. private communication). However, to our knowledge, no cleaning was applied.
Third, we visually inspected all the individual spectra to identify defects or anomalies. Although the general quality of the spectra is excellent, in a few cases, substantial anomalies were found (see Table 1 and Figure 1). The spectra were flagged for the following anomalies:
@@
Improper flux calibration at the edges and near the dichroic. These spectra were kept in the clean libraries. However, we recommend avoiding or masking the regions near the edge and near the dichroic when performing the fits.
@@
Some spectra (most commonly those coming from the Indo-US library) have localized gaps that can extend for 100s or 1000s of Å. We did not exclude them from the clean library, but we recommend excluding them if the gaps overlap with the fitted range.
@@
After the cleaning procedure, the stellar libraries consist of 789, 496, and 992 stars, as summarized in Table 1. 555The list of stars in the “clean” stellar libraries has been made available in electronic format at https://github.com/TDCOSMO/KINEMATICS_METHODS. For the clean stars in common between the libraries, we compared the spectra visually and in terms of their ratio with respect to the MILES spectrum for each star. Some examples are shown in Figure 2. Overall, the comparison is excellent, even though there are some instances in which the ratio differs significantly from one, indicating systematic residuals in the flux calibration and reddening corrections. These differences underscore the importance of having additive and multiplicative polynomial continuum components in the kinematic fits, as discussed below.
@@
| After library authors’ cutsaaaaQuality cuts applied by the authors of the libraries in the construction of SSP models (Vazdekis16; Verro22) | 801 | 530 | 1273 |
| With and Fe/HbbbbStars that do not have consistent estimates of and Fe/H (usually poor-quality spectra) | 790 | 522 | 1006 |
| After a visual inspection | 789 | 496 | 992 |
2.3 Robustness to template distribution biases
To obtain an unbiased estimate of the stellar velocity dispersion, it is crucial to use stellar libraries that span the expected temperature and metallicity ranges. We first verified that our quality-control cuts on the stellar libraries do not alter the parameter distribution, ensuring our sample remains representative.
A potential concern is that the distribution of template parameters within a library could bias the spectral fit, by acting as a prior on the recovered parameters. In this section, we show that pPXF is not affected by this issue. To test the robustness of our method, we designed an extreme experiment in which we intentionally overpopulated specific regions of the parameter space, that is, the population parameters of the spectral library.
We first constructed a mock spectrum by normalizing all E-MILES SSP templates to a unit mean flux, then co-adding different templates with weights that depend on the template age. The weights follow a Gaussian distribution centered at with , where Age is givwn in years, and we adopted an intrinsic dispersion of 100 km/s.
For clarity, we restricted the templates to solar metallicity, using the E-MILES library of stellar population models (Vazdekis16) and created a logarithmically spaced grid of stellar ages. We then generated two biased template libraries by artificially increasing the density of templates by a factor of 10, either for ages younger or older than the logarithmic midpoint (defined as , with Age in years), using linear interpolation in . Each of these biased libraries was then used with pPXF to fit the mock spectrum, using four additive polynomials and without any regularization. For each set of templates, we performed 100 random realizations of the noise to average out the noise in the recovered star formation history (SFH).
As shown in Figure 3, the averaged recovered star formation histories were statistically indistinguishable for both biased libraries. Although this experiment does not directly test the kinematics, the insensitivity of the recovered star formation history to these extreme variations provides a strong demonstration of the method’s robustness. This result implies that as long as the relevant parameter space (i.e., the population of the spectral library) is adequately covered, the fit will consistently recover the same best-fitting template and, by extension, the same kinematics, regardless of the input template distribution.